# A Survey on 3D Gaussian Splatting

### v7

@@ 3.1 Rendering with Learned 3D Gaussians
$\bullet$ Splatting. In this step, 3D Gaussians (ellipsoids) in 3D space are projected into 2D image space (ellipses). The projection proceeds through two transformations: first, transforming 3D Gaussians from world coordinates to camera coordinates using the viewing transformation, and subsequently splatting these Gaussians into 2D image space via an approximation of the projective transformation. Mathematically, given the 3D covariance matrix $\bm{\Sigma}$ describing a 3D Gaussian’s spatial distribution, and the viewing transformation matrix $\bm{W}$, the 2D covariance matrix $\bm{\Sigma}^{\prime}$ characterizing the projected 2D Gaussian is computed through:
$\bm{\Sigma}^{\prime}=\bm{JW\Sigma}\bm{W}^{\top}\bm{J}^{\top},$ (4)
where $\bm{J}$ is the Jacobian of the affine approximation of the projective transformation [39, 10]. One might wonder why the standard camera intrinsics based projective transformation is not used here. This is because its mappings are not affine and therefore cannot directly project $\bm{\Sigma}$. 3D GS adopts an affine one proposed in [39] which approximates the projective transformation using the first two terms (including $\bm{J}$) of the Taylor expansion (see Sec. 4.4 in [39]).
$x,y,z,\theta,\phi$
$c~\&~\sigma$
Figure 3: NeRFs vs. 3D GS. (a) NeRF samples along the ray and then queries the MLP to obtain corresponding colors and densities, which can be seen as a backward mapping (ray tracing). (b) In contrast, 3D GS projects all 3D Gaussians into the image space (i.e., splatting) and then performs parallel rendering, which can be viewed as a forward mapping (rasterization). Best viewed in color.
$C_{1}~C_{2}$$C_{3}~C_{4}$$C_{1}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{1}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{1}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{1}^{\prime}})$$C_{2}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{2}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{2}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{2}^{\prime}})$$C_{3}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{3}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{3}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{3}^{\prime}})$$C_{4}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{4}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{4}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{4}^{\prime}})$
Figure 4: An illustration of the forward process of 3D GS (see Sec. 3.1). (a) The splatting step projects 3D Gaussians into image space. (b) 3D GS divides the image into multiple non-overlapping patches, i.e., tiles. (c) 3D GS replicates the Gaussians which cover several tiles, assigning each copy an identifier, i.e., a tile ID. (d) By rendering the sorted Gaussians, we can obtain all pixels within the tile. Note that the computational workflows for pixels and tiles are independent and can be done in parallel. Best viewed in color.
@@
$C=\sum_{n=1}^{|\mathcal{N}|}c_{n}\alpha^{\prime}_{n}\prod_{j=1}^{n-1}\left(1-\alpha^{\prime}_{j}\right),$ (5)
@@
$\alpha^{\prime}_{n}=\alpha_{n}\times\text{exp}\big(-\frac{1}{2}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})^{\top}\bm{\Sigma}^{\prime-1}_{n}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})\big),$ (6)
@@
$\bullet$ Tiles (Patches). To avoid the cost computation of deriving Gaussians for each pixel, 3D GS shifts the precision from pixel-level to patch-level detail, which is inspired by tile-based rasterization [43]. Concretely, 3D GS initially divides the image into multiple non-overlapping patches (tiles). Fig. 4b provides an illustration of tiles. Each tile comprises $\times16$ pixels as suggested in [10]. 3D GS further determines which tiles intersect with these projected Gaussians. Given that a projected Gaussian may cover several tiles, a logical method involves replicating the Gaussian, assigning each copy an identifier (i.e., a tile ID) for the relevant tile.
$\bullet$ Parallel Rendering. After replication, 3D GS combines the respective tile ID with the depth value obtained from the view transformation for each Gaussian. This results in an unsorted list of bytes where the upper bits represent the tile ID and the lower bits signify depth. By doing so, the sorted list can be directly utilized for rendering (i.e., alpha compositing). Fig. 4c and Fig. 4d provide the visual demonstration of such concepts. It’s worth highlighting that rendering each tile and pixel occurs independently, making this process highly suitable for parallel computations. An additional benefit is that each tile’s pixels can access a common shared memory and maintain an uniform read sequence (Fig. 5), enabling parallel execution of alpha compositing with increased efficiency. In the official implementation of the original paper [10], the framework regards the processing of tiles and pixels as analogous to the blocks and threads, respectively, in CUDA programming architecture.
@@
3.2 Optimization of 3D Gaussian Splatting
@@
5
Figure 5: An illustration of the tile based parallel (at the pixel-level) rendering. All the pixels within a tile (Tile1 here) access the same ordered Gaussian list stored in a shared memory for rendering. As the system processes each Gaussian sequentially, every pixel in the tile evaluates the Gaussian’s contribution according to the distance (i.e., the $\exp$ term in Eq. 6). Therefore, the rendering for a tile can be completed by iterating through the list of Gaussians just once. The computation for the red Gaussian follows a similar way and is omitted here for simplicity.
3.2.1 Parameter Optimization
@@
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}},$ (7)
@@
3.2.2 Density Control
$\bullet$ Initialization. 3D GS starts with the initial set of sparse points from SfM or random initialization. Note that a good initialization is essential to convergence and reconstruction quality [44]. Afterwards, point densification and pruning are adopted to control the density of 3D Gaussians.
@@ 3.2.2 Density Control
4 3D Gaussian Splatting: Directions
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions. These are: i) 3D GS for Sparse Input [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (Sec. 4.1), ii) Memory-efficient 3D GS [56, 57, 58, 59, 60, 61, 62, 63, 64] (Sec. 4.2), iii) Photorealistic 3D GS [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] (Sec. 4.3), iv) Improved Optimization Algorithms [81, 22, 82, 77, 83, 84, 85, 86] (Sec. 4.4), v) 3D Gaussian with More Properties [87, 88, 89, 90, 91, 92, 93] (Sec. 4.5), vi) Hybrid Representation [94, 95, 96] (Sec. 4.6), and vii) New Rendering Algorithm (Sec. 4.7). While we have carefully selected several key directions, we acknowledge that it is inevitably a biased view. A more comprehensive collection is given in Github.
4.1 3D GS for Sparse Input
@@
Existing methods can be categorized into two primary groups. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency [49, 46, 51, 55]. For example, DNGaussian [49] introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input. FSGS [46] devised a Gaussian Unpooling process for initialization and also introduced depth regularization. MVSplat [51] proposed a cost volume representation so as to provide geometry cues. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that use learned priors [47, 48, 53, 97]. One approach involves synthesizing additional views through generative models, which can be seamlessly integrated into existing reconstruction pipelines [98]. However, this augmentation strategy is computationally intensive and inherently bounded by the capabilities of the used generative model. Another well-known paradigm employs feed-forward Gaussian model to directly generates the properties of a set of 3D Gaussians. This paradigm typically requires multiple views for training but can reconstruct 3D scenes with only one input image. For instance, PixelSplat [47] proposed to sample Gaussians from dense probability distributions. Splatter Image [48] introduced a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. However, as the generated pixel-aligned Gaussians are distributed nearly evenly in the space, they struggle to represent high-frequency details and smoother regions with an appropriate number of Gaussians.
@@ 4.2 Memory-efficient 3D GS
Recent research has pursued two primary directions to address memory efficiency. First, several approaches focus on reducing the number of 3D Gaussians [63, 58, 62]. These methods either employ strategic pruning of low-impact Gaussians, such as the volume-based masking [58], or represent neighboring Gaussians using the same properties stored within a “local anchor” obtained by clustering [22], hash-grid [62], etc. Second, researchers have developed methods for compressing Gaussian’s properties [61, 58, 62]. For instance, Niedermayr et al[61] compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning. HAC [62] predicted the probability of each quantized attribute using Gaussian distributions and then devise an adaptive quantization module. These directions are not mutually exclusive; instead, one framework might use a hybrid approach combining multiple strategies.
@@
4.3 Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians [10]. The visual fidelity of rendered images, including aspects such as aliasing, reflections, and artifacts, can be further optimized.
Recent research has focused on addressing three main aspects of visual quality, with aliasing being specific to 3D GS’s rendering algorithm, while reflection and blur handling represent broader challenges in 3D reconstruction. i) Aliasing. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Solutions emerged at both training and inference stages. Researchers developed training-time improvements from the sampling rate perspective and introduced schemes such as multi-scale Gaussians [67], 2D Mip filter [65], and conditioned logistic function [78]. Inference-time solutions, such as 2D scale-adaptive filtering [80], offer enhanced fidelity that can be integrated into any existing 3D GS frameworks. ii) Reflection. Achieving realistic rendering of reflective materials is a hard, long-standing problem in 3D scene reconstruction. Recent works have introduced various approaches to model reflective materials [68, 99, 73] and enable relightable Gaussian representation [23], though achieving physically accurate specular effects remains challenging. iii) Blur. While 3D GS excels on carefully curated datasets, real-world captures often suffer from blurs such as motion blur and defocus blur. Recent approaches explicitly incorporated blur modeling during training, employing techniques such as coarse-to-fine kernel optimization [74] and photometric bundle adjustment [75] to address this challenge.
@@ 4.3 Photorealistic 3D GS
4.4 Improved Optimization Algorithms
@@
Three main directions stand out for improving the optimization of 3D GS. i) Additional Regularization (e.g., frequency [84] and geometry [22, 77]). Geometry-aware approaches have been particularly successful, preserving scene structure through the incorporation of local anchor points [22], depth and surface constraints [100, 101, 102], Gaussian volumes [103], etc. ii) Optimization Procedure Enhancement [44, 101, 104]. While the original strategy of density control (Sec. 3.2.2) has proven valuable, considerable room for improvement remains. For example, GaussianPro [44] addresses the challenge of dense initialization in texture-less surfaces and large-scale scenes through an advanced Gaussian densification strategy. iii) Constraint Relaxation. Reliance on external tools/algorithms can introduce errors and cap the system’s performance potential. For instance, SfM, commonly used in the initialization process, is error-prone and struggle with complex scenes. Recent works have begun exploring COLMAP-free approaches utilizing stream continuity [81, 105], potentially enabling learning from internet-scale unposed video datasets.
@@
4.5 3D Gaussian with More Properties
Despite impressive, the properties of 3D Gaussian (Sec. 3.1) are designed to be used for novel-view synthesis only. By augmenting 3D Gaussian with additional properties, such as linguistic [87, 88, 89], semantic/instance [90, 91, 92], and spatial-temporal [93] properties, 3D GS demonstrates its considerable potential to revolutionize various domains.
Here we list several interesting applications using 3D Gaussians with specially designed properties. i) Language Embedded Scene Representation [87, 88, 89]. Due to the high computational and memory demands of current language-embedded scene representations, Shi et al[87] proposed a quantization scheme that augments 3D Gaussian with streamlined language embeddings instead of the original high-dimensional embeddings. This method also mitigated semantic ambiguity and enhanced the precision of open-vocabulary querying by smoothing out semantic features across different views, guided by uncertainty values. ii) Scene Understanding and Editing [90, 91, 92]. Feature 3DGS [90] integrated 3D GS with feature field distillation from 2D foundation models. By learning a lower-dimensional feature field and applying a lightweight convolutional decoder for upsampling, Feature 3DGS achieved faster training and rendering speeds while enabling high-quality feature field distillation, supporting applications like semantic segmentation and language-guided editing. iii) Spatiotemporal Modeling [93, 106]. To capture the complex spatial and temporal dynamics of 3D scenes, Yang et al[93] conceptualized spacetime as a unified entity and approximates the spatiotemporal volume of dynamic scenes using a collection of 4D Gaussians. The proposed 4D Gaussian representation and corresponding rendering pipeline are capable of modeling arbitrary rotations in space and time and allow for end-to-end training.
4.6 Hybrid Representation

### v8

@@ 3.1. Rendering with Learned 3D Gaussians
$\bullet$ Splatting (Fig. 3a). In this step, 3D Gaussians (ellipsoids) in 3D space are projected into 2D image space (ellipses). The projection proceeds through two transformations: first, transforming 3D Gaussians from world coordinates to camera coordinates using the viewing transformation, and subsequently splatting these Gaussians into 2D image space via an approximation of the projective transformation. Mathematically, given the 3D covariance matrix $\bm{\Sigma}$ describing a 3D Gaussian’s spatial distribution, and the viewing transformation matrix $\bm{W}$, the 2D covariance matrix $\bm{\Sigma}^{\prime}$ characterizing the projected 2D Gaussian is computed through:
(4) $\bm{\Sigma}^{\prime}=\bm{JW\Sigma}\bm{W}^{\top}\bm{J}^{\top},$
where $\bm{J}$ is the Jacobian of the affine approximation of the projective transformation (Zwicker et al., 2001a; Kerbl et al., 2023). One might wonder why the standard camera intrinsics based projective transformation is not used here. This is because its mappings are not affine and therefore cannot directly project $\bm{\Sigma}$. 3D GS adopts an affine one proposed in (Zwicker et al., 2001a) which approximates the projective transformation using the first two terms (including $\bm{J}$) of the Taylor expansion (see Sec. 4.4 in (Zwicker et al., 2001a)).
@@
(5) $C=\sum_{n=1}^{|\mathcal{N}|}c_{n}\alpha^{\prime}_{n}\prod_{j=1}^{n-1}\left(1-\alpha^{\prime}_{j}\right),$
@@
(6) $\alpha^{\prime}_{n}=\alpha_{n}\times\text{exp}\big(-\frac{1}{2}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})^{\top}\bm{\Sigma}^{\prime-1}_{n}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})\big),$
@@
$\bullet$ Tiles (Patches). To avoid the cost computation of deriving Gaussians for each pixel, 3D GS shifts the precision from pixel-level to patch-level detail, which is inspired by tile-based rasterization (Lassner and Zollhofer, 2021). Concretely, 3D GS initially divides the image into multiple non-overlapping patches (tiles). Fig. 3b provides an illustration of tiles. Each tile comprises $\times16$ pixels as suggested in (Kerbl et al., 2023). 3D GS further determines which tiles intersect with these projected Gaussians. Given that a projected Gaussian may cover several tiles, a logical method involves replicating the Gaussian, assigning each copy an identifier (i.e., a tile ID) for the relevant tile.
5
Figure 5. An illustration of the tile based parallel (at the pixel-level) rendering. All the pixels within a tile (Tile1 here) access the same ordered Gaussian list stored in a shared memory for rendering. As the system processes each Gaussian sequentially, every pixel in the tile evaluates the Gaussian’s contribution according to the distance (i.e., the $\exp$ term in Eq. 6). Therefore, the rendering for a tile can be completed by iterating through the list of Gaussians just once. The computation for the red Gaussian follows a similar way and is omitted here for simplicity.
$\bullet$ Parallel Rendering. After replication, 3D GS combines the respective tile ID with the depth value obtained from the view transformation for each Gaussian. This results in an unsorted list of bytes where the upper bits represent the tile ID and the lower bits signify depth. By doing so, the sorted list can be directly utilized for rendering (i.e., alpha compositing). Fig. 3c and Fig. 3d provide the visual demonstration of such concepts. It’s worth highlighting that rendering each tile and pixel occurs independently, making this process highly suitable for parallel computations. An additional benefit is that each tile’s pixels can access a common shared memory and maintain an uniform read sequence (Fig. 5), enabling parallel execution of alpha compositing with increased efficiency. In the official implementation of the original paper (Kerbl et al., 2023), the framework regards the processing of tiles and pixels as analogous to the blocks and threads, respectively, in CUDA programming architecture.
@@
3.2. Optimization of 3D Gaussian Splatting
@@
3.2.1. Parameter Optimization
@@
(7) $\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}},$
@@
3.2.2. Density Control
$\bullet$ Initialization. 3D GS starts with the initial set of sparse points from SfM or random initialization. Note that a good initialization is essential to convergence and reconstruction quality (Cheng et al., 2024). Afterwards, point densification and pruning are adopted to control the density of 3D Gaussians.
@@ 3.2.2. Density Control
4. 3D Gaussian Splatting: Directions
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions. These are: i) 3D GS for Sparse Input (Xiong et al., 2023; Zhu et al., 2024a; Charatan et al., 2024; Szymanowicz et al., 2024b; Li et al., 2024i; Swann et al., 2024; Chen et al., 2024e; Wewer et al., 2024; Xu et al., 2024b; Shen et al., 2024; Zhang et al., 2024f) (Sec. 4.1), ii) Memory-efficient 3D GS (Fan et al., 2023; Navaneet et al., 2023; Lee et al., 2024c; Morgenstern et al., 2023; Zhang et al., 2024d; Niedermayr et al., 2024; Chen et al., 2024c; Papantonakis et al., 2024; Fang and Wang, 2024) (Sec. 4.2), iii) Photorealistic 3D GS (Yu et al., 2024a; Gao et al., 2023; Yan et al., 2024b; Jiang et al., 2024b; Lee et al., 2024b; Malarz et al., 2023; Bolanos et al., 2024; Radl et al., 2024; Yang et al., 2024a; Peng et al., 2024b; Zhao et al., 2024b; Dahmani et al., 2024; Li et al., 2024f; Liang et al., 2024b; Seiskari et al., 2024; Song et al., 2024) (Sec. 4.3), iv) Improved Optimization Algorithms (Fu et al., 2024; Lu et al., 2024b; Jung et al., 2024; Li et al., 2024f; Yu et al., 2024c; Zhang et al., 2024i; Huang et al., 2024a; Li et al., 2024c) (Sec. 4.4), v) 3D Gaussian with More Properties (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024; Zhou et al., 2024a; Ye et al., 2024; Cen et al., 2023; Yang et al., 2024d) (Sec. 4.5), vi) Hybrid Representation (Yang et al., 2024b; Wu et al., 2024c; Xu et al., 2024a) (Sec. 4.6), and vii) New Rendering Algorithm (Sec. 4.7). While we have carefully selected several key directions, we acknowledge that it is inevitably a biased view. A more comprehensive collection is given in Github.
4.1. 3D GS for Sparse Input
@@
Existing methods can be categorized into two primary groups. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency (Li et al., 2024i; Zhu et al., 2024a; Chen et al., 2024e; Zhang et al., 2024f). For example, DNGaussian (Li et al., 2024i) introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input. FSGS (Zhu et al., 2024a) devised a Gaussian Unpooling process for initialization and also introduced depth regularization. MVSplat (Chen et al., 2024e) proposed a cost volume representation so as to provide geometry cues. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that use learned priors (Charatan et al., 2024; Szymanowicz et al., 2024b; Xu et al., 2024b; Szymanowicz et al., 2024a). One approach involves synthesizing additional views through generative models, which can be seamlessly integrated into existing reconstruction pipelines (Sargent et al., 2024). However, this augmentation strategy is computationally intensive and inherently bounded by the capabilities of the used generative model. Another well-known paradigm employs feed-forward Gaussian model to directly generates the properties of a set of 3D Gaussians. This paradigm typically requires multiple views for training but can reconstruct 3D scenes with only one input image. For instance, PixelSplat (Charatan et al., 2024) proposed to sample Gaussians from dense probability distributions. Splatter Image (Szymanowicz et al., 2024b) introduced a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. However, as the generated pixel-aligned Gaussians are distributed nearly evenly in the space, they struggle to represent high-frequency details and smoother regions with an appropriate number of Gaussians.
@@ 4.2. Memory-efficient 3D GS
Recent research has pursued two primary directions to address memory efficiency. First, several approaches focus on reducing the number of 3D Gaussians (Papantonakis et al., 2024; Lee et al., 2024c; Chen et al., 2024c). These methods either employ strategic pruning of low-impact Gaussians, such as the volume-based masking (Lee et al., 2024c), or represent neighboring Gaussians using the same properties stored within a “local anchor” obtained by clustering (Lu et al., 2024b), hash-grid (Chen et al., 2024c), etc. Second, researchers have developed methods for compressing Gaussian’s properties (Niedermayr et al., 2024; Lee et al., 2024c; Chen et al., 2024c). For instance, Niedermayr et al. (Niedermayr et al., 2024) compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning. HAC (Chen et al., 2024c) predicted the probability of each quantized attribute using Gaussian distributions and then devise an adaptive quantization module. These directions are not mutually exclusive; instead, one framework might use a hybrid approach combining multiple strategies.
@@
4.3. Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians (Kerbl et al., 2023). The visual fidelity of rendered images, including aspects such as aliasing, reflections, and artifacts, can be further optimized.
Recent research has focused on addressing three main aspects of visual quality, with aliasing being specific to 3D GS’s rendering algorithm, while reflection and blur handling represent broader challenges in 3D reconstruction. i) Aliasing. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Solutions emerged at both training and inference stages. Researchers developed training-time improvements from the sampling rate perspective and introduced schemes such as multi-scale Gaussians (Yan et al., 2024b), 2D Mip filter (Yu et al., 2024a), and conditioned logistic function (Liang et al., 2024b). Inference-time solutions, such as 2D scale-adaptive filtering (Song et al., 2024), offer enhanced fidelity that can be integrated into any existing 3D GS frameworks. ii) Reflection. Achieving realistic rendering of reflective materials is a hard, long-standing problem in 3D scene reconstruction. Recent works have introduced various approaches to model reflective materials (Jiang et al., 2024b; Meng et al., 2024; Yang et al., 2024a) and enable relightable Gaussian representation (Saito et al., 2024), though achieving physically accurate specular effects remains challenging. iii) Blur. While 3D GS excels on carefully curated datasets, real-world captures often suffer from blurs such as motion blur and defocus blur. Recent approaches explicitly incorporated blur modeling during training, employing techniques such as coarse-to-fine kernel optimization (Peng et al., 2024b) and photometric bundle adjustment (Zhao et al., 2024b) to address this challenge.
@@ 4.3. Photorealistic 3D GS
4.4. Improved Optimization Algorithms
@@
Three main directions stand out for improving the optimization of 3D GS. i) Additional Regularization (e.g., frequency (Zhang et al., 2024i) and geometry (Lu et al., 2024b; Li et al., 2024f)). Geometry-aware approaches have been particularly successful, preserving scene structure through the incorporation of local anchor points (Lu et al., 2024b), depth and surface constraints (Chen et al., 2023a; Yu et al., 2024d; Zhang et al., 2024c), Gaussian volumes (Chen et al., 2024d), etc. ii) Optimization Procedure Enhancement (Cheng et al., 2024; Yu et al., 2024d; Ververas et al., 2024). While the original strategy of density control (Sec. 3.2.2) has proven valuable, considerable room for improvement remains. For example, GaussianPro (Cheng et al., 2024) addresses the challenge of dense initialization in texture-less surfaces and large-scale scenes through an advanced Gaussian densification strategy. iii) Constraint Relaxation. Reliance on external tools/algorithms can introduce errors and cap the system’s performance potential. For instance, SfM, commonly used in the initialization process, is error-prone and struggle with complex scenes. Recent works have begun exploring COLMAP-free approaches utilizing stream continuity (Fu et al., 2024; Smith et al., 2024), potentially enabling learning from internet-scale unposed video datasets.
@@
4.5. 3D Gaussian with More Properties
Despite impressive, the properties of 3D Gaussian (Sec. 3.1) are designed to be used for novel-view synthesis only. By augmenting 3D Gaussian with additional properties, such as linguistic (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024), semantic/instance (Zhou et al., 2024a; Ye et al., 2024; Cen et al., 2023), and spatial-temporal (Yang et al., 2024d) properties, 3D GS demonstrates its considerable potential to revolutionize various domains.
Here we list several interesting applications using 3D Gaussians with specially designed properties. i) Language Embedded Scene Representation (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024). Due to the high computational and memory demands of current language-embedded scene representations, Shi et al. (Shi et al., 2024) proposed a quantization scheme that augments 3D Gaussian with streamlined language embeddings instead of the original high-dimensional embeddings. This method also mitigated semantic ambiguity and enhanced the precision of open-vocabulary querying by smoothing out semantic features across different views, guided by uncertainty values. ii) Scene Understanding and Editing (Zhou et al., 2024a; Ye et al., 2024; Cen et al., 2023). Feature 3DGS (Zhou et al., 2024a) integrated 3D GS with feature field distillation from 2D foundation models. By learning a lower-dimensional feature field and applying a lightweight convolutional decoder for upsampling, Feature 3DGS achieved faster training and rendering speeds while enabling high-quality feature field distillation, supporting applications like semantic segmentation and language-guided editing. iii) Spatiotemporal Modeling (Yang et al., 2024d; Lin et al., 2024a). To capture the complex spatial and temporal dynamics of 3D scenes, Yang et al. (Yang et al., 2024d) conceptualized spacetime as a unified entity and approximates the spatiotemporal volume of dynamic scenes using a collection of 4D Gaussians. The proposed 4D Gaussian representation and corresponding rendering pipeline are capable of modeling arbitrary rotations in space and time and allow for end-to-end training.
4.6. Hybrid Representation